OpenStudy (anonymous):

How many solutions are there? 3y-2x = 4 and 4x -6y = -8 ? Interpret the answer as well.

OpenStudy (anonymous):

2 solutions a x and a y

OpenStudy (anonymous):

^ Nope, wrong, when I solved I got no solution. Did you solve?

OpenStudy (anonymous):

i dont think there wld b a solution to dese pair of eqns.

OpenStudy (anonymous):

@Zeerak ...u would hv solved wrong....lol

OpenStudy (anonymous):

@fortheloveofscience ^So what should I write?

OpenStudy (anonymous):

And what does it mean when we dont have a point of intersection? What does the line represent?

OpenStudy (anonymous):

i think you should write no unique solution exists for the pair of equations

OpenStudy (anonymous):

a pair of parallel lines..

OpenStudy (anonymous):

@fortheloveofscience Thanks, and if we observe, the 2nd equation is just the multiple of the 1st, (x2) with opposite signs, what does it mean?

OpenStudy (anonymous):

They're the same equation really. For any (x,(2x+4)/3) in R^2 the equations hold.

OpenStudy (anonymous):

oh my gawsh...crap sorry its a pair of coincident lines...

OpenStudy (anonymous):

my god..how stupid i can be phew

OpenStudy (anonymous):

lol they are the same line, so infinite number of solutions

OpenStudy (anonymous):

yeah they are a pair of coincident lines so they have similar solutions...and that too infinite

OpenStudy (anonymous):

Infinite number of solutions? What does it mean? I didnt get a single pair of solution when i solved them!

OpenStudy (anonymous):

try drawing a graph..ull find urself

OpenStudy (anonymous):

No i want to knw in term of solving them algebraically, not graphing

OpenStudy (anonymous):

infinite because they overlap at every single point because they are the same graph o.o

OpenStudy (anonymous):

okay. I agree to that. But when I solved them I didnt even get a single solution, so how do we conclude algebraically that they have infinite solutions?

OpenStudy (anonymous):

okay. I agree to that. But when I solved them algebraically I didnt even get a single solution, so how do we conclude algebraically that they have infinite solutions?

OpenStudy (anonymous):

you dont have to solve them because they have the same equation

OpenStudy (anonymous):

Eq. 2 => 4x -6y = -8 => 2x - 3y = -4 Adding Eq.1 and 2: -2x + 3y = 4 2x + 3y = -4 ------------------ 0=0 They are not same equations, they have opposite signs.

OpenStudy (anonymous):

When you solve two equations, you actually assume one to be true and try to show that the other one is true too. In this case you should find some identity relation. This is a tautology, since for any number, identity is trivially true. So what you assumed should be true, that is, every (x,y) solving the equation you assumed must hold the system of equations.

OpenStudy (anonymous):

@IsaacDian Haha, whatever you wrote might be correct. But honestly it was Chinese to me. :D Don't worry I'll ask my teacher tomorrow the reason. but you get the medal for putting effort to write the answer, which definitely seems geeky. Thanks.